# Scalable and Accurate Graph Reasoning with LLM-based Multi-Agents

### v1

@@ 1 Introduction
Graphs, as a crucial data structure for modeling complex real-world relationships, are ubiquitous across various scenarios, e.g. citation networks, recommendation networks. Many important applications like drug discovery (Stokes et al., 2020), traffic forecasting (Jiang & Luo, 2022), and financial detection (Motie & Raahemi, 2024), require reasoning over graphs to be realized. Noticing the powerful general knowledge and language processing capabilities of Large Language Models (LLMs) (Brown et al., 2020), a significant amount of works have focused on using LLMs to perform various reasoning tasks, such as mathematical formula derivation (Meadows et al., 2023), commonsense reasoning (Madaan et al., 2022), and multi-hop question answering (Creswell et al., 2023). However, most of them primarily involve shallow or sequential reasoning. To bring the LLM reasoning closer to human thinking, it is necessary for LLMs to master deeper and more complex reasoning, such as graph reasoning.
@@
Despite significant efforts by researchers to enable LLMs to memorize, comprehend, and perform basic reasoning on graph structures, several issues still persist: 1) The scale of graphs that can be handled is limited. Describing graph structures in natural language inevitably leads to excessively long inputs. Due to context length limitations and the shortcomings of LLMs in handling lengthy text (Liu et al., 2023), previous works (Chai et al., 2023; Fatemi et al., 2024; Perozzi et al., 2024) could only handle graphs of very limited size (e.g. fewer than 20 nodes and 100 edges). 2) The performance on graph reasoning tasks is relatively poor. Unlike text, which can tolerate some degree of semantic deviation, reasoning and computation on graphs must be highly precise. However, current works demonstrate poor accuracy (average 2060%) in various graph reasoning tasks like connectivity and shortest path. 3) Lacking explicit reasoning paths. Taking the shortest path as an example, the responses of existing models resemble a heuristic search approach to finding the shortest path on a graph, rather than strictly executing an algorithm. This makes it difficult to determine whether LLMs are genuinely deriving the answer through correct reasoning or merely making educated guesses. Although GraphWiz (Chen et al., 2024a) attempts to generate explicit reasoning paths through fine-tuning, it often fails due to the presence of incomplete or wrong reasoning paths in its training data. Furthermore, GraphWiz exhibits overfitting, where it tends to treat new or unrelated questions as one of the fine-tuned problems, which will be detailed in Section 5.3.
@@
Motivation. The ultimate goal of graph reasoning is to enable LLMs to leverage graph-related knowledge or algorithms to solve real-world graph problems. However, with the development of information science and hardware storage, the scale of graphs and information per node become too large for a single LLM to handle. To address this, a natural idea is to use distributed approaches, where a large graph is stored across multiple LLMs separately and compute collaboratively. Therefore, just as graph algorithms have generally evolved from non-distributed to distributed forms (Meng et al., 2024b)), we hope that LLMs can also learn the concept of distributed processing, thereby harnessing the power of swarm intelligence to solve graph problems in real-world scenarios.
@@ 2 Preliminaries and Related Works
Large Language Models for Graph Reasoning. To further enhance the reasoning capabilities of LLMs, many works have attempted to improve the performance of LLMs in graph reasoning. Wang et al. (2023) first introduces the NLGraph Benchmark to evaluate the performance of LLMs on various graph reasoning tasks. Fatemi et al. (2024) explores the impact of different graph encoding methods and graph structure types on the performance of LLMs in graph reasoning tasks. Additionally, it introduces another benchmark called GraphQA. Considering the lengthy nature of describing graph structures in text, Chai et al. (2023) and Perozzi et al. (2024) respectively use Transformers and GNNs to encode graph structures and attempt to align them with LLMs. Inspired by how humans understand structural information through the visual modality, Wei et al. (2024) generates corresponding visual images based on graph structures and provides them to visual LLMs for graph reasoning. Chen et al. (2024a) conducted Supervised Fine-Tuning and Directly Prefered Optimization on LLMs, enhancing the performance of LLMs and encouraging them to output explicit reasoning paths.
Large Language Model based Multi-Agents. Recent advancements in LLMs have spurred interest in their application within multi-agent systems. LLM-based multi-agent frameworks leverage the natural language understanding and reasoning capabilities of LLMs to enable agents to collaborate, communicate, and solve complex tasks in a distributed manner. Existing multi-agents works for problem solving primarily focuses on applications such as Software Development (Dong et al., 2023; Hong et al., 2024; Qian et al., 2024), Embodied Agents (Zhang et al., 2024; Mandi et al., 2024; Chen et al., 2024b) and Science Debate (Xiong et al., 2023; Chan et al., 2024). However, using LLM-based multi-agents to handle graph data has been less explored, especially in the areas of graph reasoning and graph computation tasks. This may be due to the hallucination issue inherent in LLMs (Huang et al., 2023), where their responses are factually incorrect. This problem becomes more complex in a multi-agent setting, as the hallucinations of a single agent may propagate to other nodes by communication (Guo et al., 2024). This requires the performance of individual agents be sufficiently stable to ensure the correct operation of the entire multi-agent system.
@@ 3 Limitations of Single LLM in Graph Reasoning
Although LLMs exhibit strong language processing and logical reasoning capabilities, problems with the Transformer architecture and Attention mechanism (Vaswani et al., 2017) still limit the scale and accuracy when they process graph problems. There are two primary limitations:
@@
The graph structure is too complex to memorize and understand for a single LLM. Using adjacency or edge lists to describe graph structures in natural language is the most intuitive and direct method, facilitating the processing of graph data by LLMs through text. However, this approach inevitably leads to a lengthy context, as the number of edges can grow quadratically with the number of nodes. As the graph scales up and becomes denser, the graph structure becomes highly complex, requiring a large amount of tokens to describe the edge relationships. When the text becomes too lengthy, it becomes difficult for LLMs to properly allocate attention, and they may even struggle with simple tasks such as key-value pair matching Liu et al. (2023). This presents significant challenges for LLMs in identifying key information for graph reasoning tasks from the lengthy context. Figure 2 shows the performance of a single LLM in memorizing one-hop neighbor nodes. We observe that as the number of nodes in the graph increases, various LLMs exhibit a significant decline in accuracy. If a single LLM cannot even correctly recall basic graph structural information like node neighbors, it becomes difficult to proceed with more complex graph reasoning or computation.
@@
Furthermore, the graph structure is described in a sequential manner. LLMs have to identify implicit graph structures from sequential text. Since the processing of LLMs is a black-box operation, it is difficult to assert that they truly construct graph structures implicitly and thereby understand them. Huang et al. (2024) conducted extensive experiments to explore whether LLMs treat the input prompts as graphs or merely as paragraphs with keywords on TAGs. The results show that the performance of LLMs in handling TAGs primarily stems from the context rather than the graph structure. LLMs tend to process the graph description as linearized paragraphs rather than graphs.
@@
A single LLM struggles to solve reasoning problems in real-world scenarios. Researchers train LLMs on graph reasoning tasks to empower them to utilize learned graph-related knowledge or algorithms to tackle real-world graph problems. However, in practical scenarios, the amount of information associated with each node can be enormous. Take citation networks as an example: a single node represents a paper, and its node information includes the title, abstract, and references, which could amount to several thousand tokens. In addition to the complexity of graph structures, the need to handle a large amount of node information further exacerbates the burden on a single LLM and highlights its shortcomings in processing long contexts. Moreover, using a single LLM to handle the entire network is inefficient, as it cannot coherently process the entire network’s problems. Typically, it is necessary to manually compress or summarize the information for each node and then feed local subgraphs to the LLM for processing (Guo et al., 2023; Chen et al., 2023).
@@
Furthermore, many current works (Chen et al., 2024a; Perozzi et al., 2024) require training GNNs or fine-tuning LLMs on individual or multiple graph reasoning tasks. However, when transferring to other graph tasks, a certain degree of performance degradation occurs, and retraining or fine-tuning for new graph tasks consumes a significant amount of time and resources. Whether LLMs can apply the graph knowledge and algorithms learned during the training process to actual graph reasoning also remains an open question. We explored this question in 5.3 and observed significant overfitting in LLMs fine-tuned on specific graph reasoning tasks. Therefore, the ideal solution would be to leverage the powerful general knowledge acquired during the pre-training phase of LLMs through an appropriate approach, enabling them to handle graph reasoning tasks as naturally as they do with natural language problems.
@@ 4 GraphAgent-Reasoner
To solve the limitations above, we propose a novel framework based on multi-agent collaboration called GraphAgent-Reasoner as shown in Figure 3, aiming to solve graph reasoning problems explicitly and correctly. The interface of the framework is a Master LLM, which is responsible for processing the textual input of graph problems, constructing the agent network, directing them to collaboratively solve the problem, and finally aggregating the states of all agents to derive the solution. Its implementation is based on the React Agent proposed by Yao et al. (2023), which is capable of reasoning based on the environment and executing corresponding actions, as detailed later. The pipeline of GAR consists of four steps: Graph Construction, Algorithm Establishing, Distributed Execution and Master Summarization.
@@
Since LLMs lack prior knowledge of this distributed paradigm, to facilitate the Master LLM’s understanding and application of the framework, we develop a distributed algorithm library that adheres to this distributed paradigm, from which the Master LLM can query relevant algorithm templates to generate distributed solutions within this paradigm. Specifically, we selected classic distributed graph algorithms and documented their implementations under this distributed paradigm. Some examples are presented in Appendix A.1. Drawing on prior work (Zheng et al., 2024; Meng et al., 2024a), we endeavor to write detailed reasoning steps of each part in the algorithm to encourage the agent to think step by step as much as possible, which plays an important role in enhancing the success rate of individual agents.
@@ 5 Experiments
Datasets. We conduct our experiments on the graph reasoning tasks proposed in GraphInstruct (Chen et al., 2024a). This dataset contains nine graph reasoning problems with different time complexity, ranging from linear and polynomial complexity to NP-complete.
@@
Setting. The underlying reasoning LLM of Agent Node used in our framework is ChatGPT-4o-mini-2024-07-18, and the base model of Master LLM is ChatGPT-4-turbo(OpenAI, 2023). The temperature is consistently set to 0. Our framwork is built upon AgentScope (Gao et al. (2024)), an innovative platform to easily build reliable, high-performance multi-agent applications.
@@ 5.2 Experiment 2: Performance on Large-Scale Graphs
In this experiment, we evaluate the performance of current LLMs on large-scale graphs. The largest graph size handled by existing graph reasoning work is 100 nodes (Chen et al., 2024a), which is still far from sufficient for real-world graph reasoning scenarios. To evaluate the reasoning performance of existing models on larger graphs, we conduct shortest path experiments on graphs with 100, 200, 500, and 1000 nodes. Due to the excessively long input text (reaching 16,000 tokens for 1000 nodes) and the money cost, we only create 20 test samples for each graph size. The results are shown in Table 2.
@@
We see the two GraphWiz models fine-tuned on the LLaMA2-7B (Touvron et al., 2023) base model are unable to handle graphs with 500 or more nodes due to the context length limitation (the context length limit for Llama2 is 4096 tokens). Although ChatGPT-3.5-turbo-16k and ChatGPT-4-32k can manage longer contexts, they output wrong answers in almost all test samples, with only ChatGPT-4-32k being correct in one 200 nodes test sample. In contrast, GAR maintains a high accuracy in large-scale graph, only failed in two 1000-node test samples, further demonstrating its robust scalability.
@@ 5.3 Experiment 3: Case Study
Although GraphWiz performed well on fine-tuned tasks, it exhibits severe overfitting when faced with real-world graph problems, failing to apply the graph reasoning knowledge learned during the fine-tuning phase. Since GraphWiz uses a consistent graph node description, the sentence "The nodes are numbered from 0 to …" appears across all datasets during the mixed-task instruction tuning. When the actual problem has nodes numbered from 1 to 20, it still assumes the existence of node 0. As a result, both GraphWiz models first output that the graph has 21 nodes and an incorrect number of edges. Furthermore, neither of the two GraphWiz models recognizes that this is a problem associated with web page importance ranking. Instead, they approach it as the bipartite graph check or topological sort problems they had been fine-tuned on. Additionally, neither model generates an explicit and correct reasoning path. These observations indicate that there is still a significant gap between excelling in classic graph reasoning tasks and effectively solving real-world graph reasoning problems. In contrast, GAR correctly identifies that the problem should be solved using knowledge related to PageRank (Yang et al., 2024) and designs an algorithm that adhered to the distributed paradigm (Note: the distributed algorithm library does not contain a PageRank algorithm template). GAR then assigns the algorithm to agent nodes for execution, ultimately obtaining the PageRank value for each node and arriving at the correct conclusion. Through the distributed paradigm, GAR effectively bridges the powerful knowledge learned by LLMs with the solving of real-world graph reasoning problems, which enables it to flexibly handle practical issues in a distributed manner. This case study demonstrates the feasibility of using GAR to solve real-world graph reasoning problems, indicating its substantial practical applicability and offering researchers and practitioners a powerful framework to address such tasks.

### v2

@@ 1 Introduction
Graphs, as a crucial data structure for modeling complex real-world relationships, are ubiquitous across various scenarios, e.g. citation networks, recommendation networks. Many important applications like drug discovery (stokes2020deep), traffic forecasting (jiang2022graph), and financial detection (motie2023financial), require reasoning over graphs to be realized. Noticing the powerful general knowledge and language processing capabilities of Large Language Models (LLMs) (brown2020language), a significant amount of works have focused on using LLMs to perform various reasoning tasks, such as mathematical formula derivation (meadows2023math), commonsense reasoning (Madaan2022commonsense), and multi-hop question answering (Creswell2023selection). However, most of them primarily involve shallow or sequential reasoning. To bring the LLM reasoning closer to human thinking, it is necessary for LLMs to master deeper and more complex reasoning, such as graph reasoning.
@@
Despite significant efforts by researchers to enable LLMs to memorize, comprehend, and perform basic reasoning on graph structures, several issues still persist: 1) The scale of graphs that can be handled is limited. Describing graph structures in natural language inevitably leads to excessively long inputs. Due to context length limitations and the shortcomings of LLMs in handling lengthy text (Liu2023LostIT), previous works (chai2023graphllm; fatemi2023talk; perozzi2024let) could only handle graphs of very limited size (e.g. fewer than 20 nodes and 100 edges). 2) The performance on graph reasoning tasks is relatively poor. Unlike text, which can tolerate some degree of semantic deviation, reasoning and computation on graphs must be highly precise. However, current works demonstrate poor accuracy (average 2060%) in various graph reasoning tasks like connectivity and shortest path. 3) Lacking explicit reasoning paths. Taking the shortest path as an example, the responses of existing models resemble a heuristic search approach to finding the shortest path on a graph, rather than strictly executing an algorithm. This makes it difficult to determine whether LLMs are genuinely deriving the answer through correct reasoning or merely making educated guesses. Although GraphWiz (chen2024graphwiz) attempts to generate explicit reasoning paths through fine-tuning, it often fails due to the presence of incomplete or wrong reasoning paths in its training data. Furthermore, GraphWiz exhibits overfitting, where it tends to treat new or unrelated questions as one of the fine-tuned problems, which will be detailed in Section 5.3.
@@
Motivation. The ultimate goal of graph reasoning is to enable LLMs to leverage graph-related knowledge or algorithms to solve real-world graph problems. However, with the development of information science and hardware storage, the scale of graphs and information per node become too large for a single LLM to handle. To address this, a natural idea is to use distributed approaches, where a large graph is stored across multiple LLMs separately and compute collaboratively. Therefore, just as graph algorithms have generally evolved from non-distributed to distributed forms (meng2024survey)), we hope that LLMs can also learn the concept of distributed processing, thereby harnessing the power of swarm intelligence to solve graph problems in real-world scenarios.
@@ 2 Preliminaries and Related Works
Large Language Models for Graph Reasoning. To further enhance the reasoning capabilities of LLMs, many works have attempted to improve the performance of LLMs in graph reasoning. wang2024can first introduces the NLGraph Benchmark to evaluate the performance of LLMs on various graph reasoning tasks. fatemi2023talk explores the impact of different graph encoding methods and graph structure types on the performance of LLMs in graph reasoning tasks. Additionally, it introduces another benchmark called GraphQA. Considering the lengthy nature of describing graph structures in text, chai2023graphllm and perozzi2024let respectively use Transformers and GNNs to encode graph structures and attempt to align them with LLMs. Inspired by how humans understand structural information through the visual modality, Wei2024GITAGT generates corresponding visual images based on graph structures and provides them to visual LLMs for graph reasoning. chen2024graphwiz conducted Supervised Fine-Tuning and Directly Prefered Optimization on LLMs, enhancing the performance of LLMs and encouraging them to output explicit reasoning paths.
Large Language Model based Multi-Agents. Recent advancements in LLMs have spurred interest in their application within multi-agent systems. LLM-based multi-agent frameworks leverage the natural language understanding and reasoning capabilities of LLMs to enable agents to collaborate, communicate, and solve complex tasks in a distributed manner. Existing multi-agents works for problem solving primarily focuses on applications such as Software Development (dong2023self; hong2024metagpt; qian2024chatdev), Embodied Agents (zhang2024build; zhao2024roco; chen2024scalable) and Science Debate (xiong2023examing; chan2024chateval). However, using LLM-based multi-agents to handle graph data has been less explored, especially in the areas of graph reasoning and graph computation tasks. This may be due to the hallucination issue inherent in LLMs (huang2023hallucination), where their responses are factually incorrect. This problem becomes more complex in a multi-agent setting, as the hallucinations of a single agent may propagate to other nodes by communication (guo2024multiagent). This requires the performance of individual agents be sufficiently stable to ensure the correct operation of the entire multi-agent system.
@@ 3 Limitations of Single LLM in Graph Reasoning
Although LLMs exhibit strong language processing and logical reasoning capabilities, problems with the Transformer architecture and Attention mechanism (vaswani2017attention) still limit the scale and accuracy when they process graph problems. There are two primary limitations:
@@
The graph structure is too complex to memorize and understand for a single LLM. Using adjacency or edge lists to describe graph structures in natural language is the most intuitive and direct method, facilitating the processing of graph data by LLMs through text. However, this approach inevitably leads to a lengthy context, as the number of edges can grow quadratically with the number of nodes. As the graph scales up and becomes denser, the graph structure becomes highly complex, requiring a large amount of tokens to describe the edge relationships. When the text becomes too lengthy, it becomes difficult for LLMs to properly allocate attention, and they may even struggle with simple tasks such as key-value pair matching Liu2023LostIT. This presents significant challenges for LLMs in identifying key information for graph reasoning tasks from the lengthy context. Figure 2 shows the performance of a single LLM in memorizing one-hop neighbor nodes. We observe that as the number of nodes in the graph increases, various LLMs exhibit a significant decline in accuracy. If a single LLM cannot even correctly recall basic graph structural information like node neighbors, it becomes difficult to proceed with more complex graph reasoning or computation.
@@
Furthermore, the graph structure is described in a sequential manner. LLMs have to identify implicit graph structures from sequential text. Since the processing of LLMs is a black-box operation, it is difficult to assert that they truly construct graph structures implicitly and thereby understand them. huang2024can conducted extensive experiments to explore whether LLMs treat the input prompts as graphs or merely as paragraphs with keywords on TAGs. The results show that the performance of LLMs in handling TAGs primarily stems from the context rather than the graph structure. LLMs tend to process the graph description as linearized paragraphs rather than graphs.
@@
A single LLM struggles to solve reasoning problems in real-world scenarios. Researchers train LLMs on graph reasoning tasks to empower them to utilize learned graph-related knowledge or algorithms to tackle real-world graph problems. However, in practical scenarios, the amount of information associated with each node can be enormous. Take citation networks as an example: a single node represents a paper, and its node information includes the title, abstract, and references, which could amount to several thousand tokens. In addition to the complexity of graph structures, the need to handle a large amount of node information further exacerbates the burden on a single LLM and highlights its shortcomings in processing long contexts. Moreover, using a single LLM to handle the entire network is inefficient, as it cannot coherently process the entire network’s problems. Typically, it is necessary to manually compress or summarize the information for each node and then feed local subgraphs to the LLM for processing (guo2023gpt4graph; chen2024exploring).
@@
Furthermore, many current works (chen2024graphwiz; perozzi2024let) require training GNNs or fine-tuning LLMs on individual or multiple graph reasoning tasks. However, when transferring to other graph tasks, a certain degree of performance degradation occurs, and retraining or fine-tuning for new graph tasks consumes a significant amount of time and resources. Whether LLMs can apply the graph knowledge and algorithms learned during the training process to actual graph reasoning also remains an open question. We explored this question in 5.3 and observed significant overfitting in LLMs fine-tuned on specific graph reasoning tasks. Therefore, the ideal solution would be to leverage the powerful general knowledge acquired during the pre-training phase of LLMs through an appropriate approach, enabling them to handle graph reasoning tasks as naturally as they do with natural language problems.
@@ 4 GraphAgent-Reasoner
To solve the limitations above, we propose a novel framework based on multi-agent collaboration called GraphAgent-Reasoner as shown in Figure 3, aiming to solve graph reasoning problems explicitly and correctly. The interface of the framework is a Master LLM, which is responsible for processing the textual input of graph problems, constructing the agent network, directing them to collaboratively solve the problem, and finally aggregating the states of all agents to derive the solution. Its implementation is based on the React Agent proposed by yao2023react, which is capable of reasoning based on the environment and executing corresponding actions, as detailed later. The pipeline of GAR consists of four steps: Graph Construction, Algorithm Establishing, Distributed Execution and Master Summarization.
@@
Since LLMs lack prior knowledge of this distributed paradigm, to facilitate the Master LLM’s understanding and application of the framework, we develop a distributed algorithm library that adheres to this distributed paradigm, from which the Master LLM can query relevant algorithm templates to generate distributed solutions within this paradigm. Specifically, we selected classic distributed graph algorithms and documented their implementations under this distributed paradigm. Some examples are presented in Appendix A.1. Drawing on prior work (zheng2024executing; jojic2024gpt), we endeavor to write detailed reasoning steps of each part in the algorithm to encourage the agent to think step by step as much as possible, which plays an important role in enhancing the success rate of individual agents.
@@ 5 Experiments
Datasets. We conduct our experiments on the graph reasoning tasks proposed in GraphInstruct (chen2024graphwiz). This dataset contains nine graph reasoning problems with different time complexity, ranging from linear and polynomial complexity to NP-complete.
@@
Setting. The underlying reasoning LLM of Agent Node used in our framework is ChatGPT-4o-mini-2024-07-18, and the base model of Master LLM is ChatGPT-4-turbo(achiam2023gpt). The temperature is consistently set to 0. Our framwork is built upon AgentScope (agentscope), an innovative platform to easily build reliable, high-performance multi-agent applications.
@@ 5.2 Experiment 2: Performance on Large-Scale Graphs
In this experiment, we evaluate the performance of current LLMs on large-scale graphs. The largest graph size handled by existing graph reasoning work is 100 nodes (chen2024graphwiz), which is still far from sufficient for real-world graph reasoning scenarios. To evaluate the reasoning performance of existing models on larger graphs, we conduct shortest path experiments on graphs with 100, 200, 500, and 1000 nodes. Due to the excessively long input text (reaching 16,000 tokens for 1000 nodes) and the money cost, we only create 20 test samples for each graph size. The results are shown in Table 2.
@@
We see the two GraphWiz models fine-tuned on the LLaMA2-7B (Touvron2023llama2) base model are unable to handle graphs with 500 or more nodes due to the context length limitation (the context length limit for Llama2 is 4096 tokens). Although ChatGPT-3.5-turbo-16k and ChatGPT-4-32k can manage longer contexts, they output wrong answers in almost all test samples, with only ChatGPT-4-32k being correct in one 200 nodes test sample. In contrast, GAR maintains a high accuracy in large-scale graph, only failed in two 1000-node test samples, further demonstrating its robust scalability.
@@ 5.3 Experiment 3: Case Study
Although GraphWiz performed well on fine-tuned tasks, it exhibits severe overfitting when faced with real-world graph problems, failing to apply the graph reasoning knowledge learned during the fine-tuning phase. Since GraphWiz uses a consistent graph node description, the sentence "The nodes are numbered from 0 to …" appears across all datasets during the mixed-task instruction tuning. When the actual problem has nodes numbered from 1 to 20, it still assumes the existence of node 0. As a result, both GraphWiz models first output that the graph has 21 nodes and an incorrect number of edges. Furthermore, neither of the two GraphWiz models recognizes that this is a problem associated with web page importance ranking. Instead, they approach it as the bipartite graph check or topological sort problems they had been fine-tuned on. Additionally, neither model generates an explicit and correct reasoning path. These observations indicate that there is still a significant gap between excelling in classic graph reasoning tasks and effectively solving real-world graph reasoning problems. In contrast, GAR correctly identifies that the problem should be solved using knowledge related to PageRank (yang2024pagerank) and designs an algorithm that adhered to the distributed paradigm (Note: the distributed algorithm library does not contain a PageRank algorithm template). GAR then assigns the algorithm to agent nodes for execution, ultimately obtaining the PageRank value for each node and arriving at the correct conclusion. Through the distributed paradigm, GAR effectively bridges the powerful knowledge learned by LLMs with the solving of real-world graph reasoning problems, which enables it to flexibly handle practical issues in a distributed manner. This case study demonstrates the feasibility of using GAR to solve real-world graph reasoning problems, indicating its substantial practical applicability and offering researchers and practitioners a powerful framework to address such tasks.